# Model Predictive Adversarial Imitation Learning for Planning from Observation

### v1

@@ Abstract
Human demonstration data is often ambiguous and incomplete, motivating imitation learning approaches that also exhibit reliable planning behavior. A common paradigm to perform planning-from-demonstration involves learning a reward function via Inverse Reinforcement Learning (IRL) then deploying this reward via Model Predictive Control (MPC). Towards unifying these methods, we derive a replacement of the policy in IRL with a planning-based agent. With connections to Adversarial Imitation Learning, this formulation enables end-to-end interactive learning of planners from observation-only demonstrations. In addition to benefits in interpretability, complexity, and safety, we study and observe significant improvements on sample efficiency, out-of-distribution generalization, and robustness. The study includes evaluations in both simulated control benchmarks and real-world navigation experiments using few-to-single observation-only demonstrations.
Keywords: Imitation Learning, Model Predictive Control, Inverse Reinforcement Learning
@@ 1 Introduction
Figure 1: Model Predictive Adversarial Imitation Learning (MPAIL) embeds an MPC agent within Adversarial Imitation Learning (AIL) for learning costs from observation. On deployment, the MPC agent performs online optimization through short model rollouts, or plans, as they are evaluated by the learned costs and value networks.
Inverse Reinforcement Learning (IRL) offers a principled approach to imitation learning by inferring the underlying reward function that explains expert behavior. A major advantage of IRL is its ability to generalize beyond demonstration data, enabling agents to generate new plans aligned with the expert’s objectives. This feature allows for improved interpretability in decision making, even when demonstrations are ambiguous, suboptimal, or complex [1]. As such, IRL is particularly well-suited for domains where understanding preferences and ensuring reliable planning are essential, such as routing on Google Maps [2], socially aware navigation [3], and autonomous driving [4].
For real-time systems, learned IRL and Inverse Optimal Control (IOC) rewards are typically deployed via Model Predictive Control (MPC) [5, 6, 7, 8, 9, 10, 11]. Here, the offline IRL algorithm iteratively solves a Reinforcement Learning (RL) problem in an inner loop, guided by the current reward estimate. An outer loop then updates this reward to minimize the discrepancy between the agent’s and the expert’s behavior. Once training is complete, the resulting reward is integrated with MPC for real-time planning and control.
Adversarial Imitation Learning (AIL) has made significant improvements over IRL in algorithmic complexity and sample efficiency [12, 13]. However, the reliance on an RL policy in AIL methods complicates their use in applications with safety constraints [5, 7, 8]. Further limited by partial observability, these deployments will often prioritize planning using a model for the sake of real-time performance, trustworthiness, and interpretability [14, 15, 16].
In this work, we propose to replace the standard policy network in AIL with an MPC agent, yielding key benefits:
@@
    Planning-from-Observation. Towards scalable Learning-from-Observation (LfO), a predictive model can be used to match the expert without requiring action data. In addition, MPC naturally admits interpretable model-based planning, which can provide insight into the agent’s decision making process as it learns from complex expert data. We further show that this can improve sample efficiency, out-of-distribution recovery, and enable navigation from a single partially observable demonstration.
@@
    Unification of IRL and MPC. These methods are otherwise considered as independent training and deployment procedures. By embedding the planner directly in the learning pipeline, online settings (e.g., dynamics, preferences, control constraints) can be brought into training and further allow for long-horizon reasoning through on-policy value function estimation. We find that this induces a more effective adversarial dynamic and enables MPC to robustly learn from observation beyond its planning horizon and data coverage.
By choosing Model Predictive Path Integral control (MPPI) [17] as the embedded planner, we derive theoretical connections to the Adversarial Imitation Learning objective. Thus, we name this learning framework, Model Predictive Adversarial Imitation Learning (MPAIL) {pronounced: impale}. We implement MPAIL and empirically validate it as a planning-based AIL algorithm. From a single real-world partially observable demonstration, MPAIL enables interpretable real-time planning from observation on limited compute through a Real-Sim-Real pipeline.
@@ 2 Related Work
IRL-MPC. High-dimensional continuous control applications often require an online planner for real-time control, trustworthiness, safety, or additional constraints. When using IRL to learn a reward, online deployments of these reward functions tend to rely on an independent online MPC procedure. To enable learning local costmaps across perception and control for off-road navigation, Lee et al. [5] and Triest et al. [7] similarly propose solving the forward RL problem by using MPPI but deploy the learned reward on a different configuration more suitable for real-time planning and control. This reward deployment framework of IRL-then-MPC is currently the dominant approach for planning in high-dimensional continuous control tasks from demonstration [5, 6, 7, 8, 9, 10, 11].
Model-Based IRL and Planning-Based RL. The proposed framework, MPAIL, can be viewed as a model-based AIL approach. Various other works have also explored model-based AIL [18, 4, 19]. However, scope is directed at training stabilization and policy performance rather than planning using the learned rewards. Notably, TD-MPC [20], a model-based RL algorithm has demonstrated considerable improvement over existing state-of-the-art RL algorithms through latent state planning. MPAIL similarly performs local trajectory optimization and learns a terminal value function for long-horizon reasoning. This work does not include a policy-based prior to better isolate and study the effects of the planning-based agent [21, 19].
@@ 3.1 The POMDP Setting and the Model Predictive Agent
We adopt the Partially Observable Markov Decision Process (POMDP) to best consider highly desirable applications of IRL in which partial observability and model-based planning play crucial roles. In an unknown world state $s_{w}\in\mathcal{S}_{w}$, the agent makes an observation $o\sim p(o|s_{w})$. From a history of observations $\mathbf{o}_{0:t}$, the Agent perceives its state $s_{t}\sim p(s|\mathbf{o}_{0:t})$. Actions $a\in\mathcal{A}$ and states $s\in\mathcal{S}$ together allow the agent to self-predict forward in time using its predictive model $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. Note that these definitions crucially suggest the partial observability of $s$ due to the implicit dependency on the observation history $\mathbf{o}_{0:t}$ through the agent’s perception (e.g. mapping [22]). However, partial observations in $\mathcal{S}$ are desirably used for demonstrations to perform IRL and AIL, as full observation history would quickly become intractable [7]. The planner itself is a model predictive agent. It is capable of performing model rollouts $\tau^{(H)}_{t}=\{s_{t^{\prime}},a_{t^{\prime}}\}_{t^{\prime}=t}^{t+H}$ such that $s_{t^{\prime}+1}=f(s_{t^{\prime}},a_{t^{\prime}})$. Each rollout is then mapped to a cost $C(\tau_{t})$. Finally, the agent’s objective is to create an $H$-step action sequence $\mathbf{a}_{t:t+H}$, or plan, that best minimizes its corresponding trajectory cost.
@@ 3.2 Adversarial Imitation Learning from Observation
IRL algorithms aim to learn a cost function that minimizes the cost of expert trajectories while maximizing the cost of trajectories induced by other policies [23, 12]. As the problem is ill-posed and many costs can correspond to a given set of demonstrations, the principle of maximum entropy is imposed to obtain a uniquely optimal cost. It can be shown that the entropy maximizing distribution is a Boltzmann distribution [1]. Towards scalable learning-from-observation (LfO), we further consider demonstration data in which only states are available, as actions can be challenging or impossible to obtain [24]. In this context, the state-only IRL from observation problem can be formulated by costing state-transitions $c(s,s^{\prime})$ rather than state-actions $c(s,a)$ as in [23]:
$\text{IRLfO}_{\psi}(\pi_{E})=\arg\max_{c\in\mathbb{R}^{\mathcal{S}\times\mathcal{S}}}-\psi(c)+\left(\min_{\pi\in\Pi}-\lambda\mathbb{H}(\pi)+\mathbb{E}_{\pi}[c(s,s^{\prime})]\right)-\mathbb{E}_{\pi_{E}}[c(s,s^{\prime})],$ (1)
@@
As shown in [12, 23], this objective can be shown to be dual to the Adversarial Imitation Learning (AIL) objective under a specific choice of cost regularizer $\psi$,
@@
where $D(\cdot)$ is the discriminator function. The exact form of $D$ has consequences on the policy objective and differs by AIL algorithm. The following section expands on choices for this work.
@@ 3.3 Choosing a Policy and Reward
In brief, we set our sights on the AIL objective (Equation 2) which, dual to MaxEnt IRL, aims to learn rewards from demonstration. However, the formulation remains intimately connected with policy optimization through the assumption of an RL procedure. Towards planning-based optimization, we proceed by modifying the RL formulation in Equation 1 and described in [12, 21]. Similar to [25], we replace the entropy loss $-\lambda\mathbb{H}(\pi)$ with a Kullback-Leibeler (KL) divergence constraint on the previous policy $\overline{\pi}$:
@@
We then observe that a choice of planner satisfies the RL objective as in Equation 3. By choosing Model Predictive Path Integral (MPPI) as the planner, as proven in Section B.1, we solve an equivalent problem provided the MDP is uniformly ergodic. Namely, MPPI solves for a KL-constrained cost-minimizer over trajectories [25]:
@@
For practical reasons, model rollouts are often limited to some timestep length, $H$. However, this can often result in myopic plans or limit applications to short-horizon tasks [26]. To resolve this, infinite-horizon MPPI introduces a terminal cost-to-go function to be evaluated on the final states in the rollouts [26, 27, 28, 29]. This can be done by utilizing a learned value function $V_{\phi}:\mathcal{S}\rightarrow\mathbb{R}$ that estimates the expected return $G_{t}$ of a state $s_{t}$ as $G_{t}=\mathbb{E}_{\pi}[R_{t+1}+\gamma R_{t+2}+\gamma^{2}R_{t+3}+...|S_{t}=s_{t}]$, where $R_{t+1}=R(s_{t},s_{t+1})$. The result of MPPI’s approximately global policy optimization at each timestep is what is referred to as the MPPI policy, $\pi_{\text{MPPI}}$. Section B.1 proves how this formulation can be equivalent to the entropy-regularized RL objective, while also in the observation-only setting. Pseudocode for the MPPI procedure can be found in Algorithm 2 as well as for its adaptation as an RL policy in Algorithm 3. Figure 2 illustrates the policy.
Figure 2: Illustration of $\pi_{\text{MPPI}}$ in MPAIL. (1) A set of action sequences (plans) are sampled and rolled out. (2) Plans are costed according to the discriminator, shifting the distribution towards the expert. Temperature $\lambda$ decreases during training, narrowing the optimized distribution in subsequent episodes. (3) The policy $\pi_{\text{MPPI}}$ is the result of a Gaussian fit to the optimized plans and their respective first actions.
In short, our chosen AIL agent, infinite-horizon MPPI, can be viewed as an approximately global optimizer for an entropy-regularized policy at a given state. Borrowing from the online learning perspective [30], these optimizations occur online rather than offline as part of, for instance, the canonical actor-critic update.
Provided the agent (policy), we now proceed with selecting its objective. Recent work has shown many potential choices of valid policy objectives, each with various empirical tradeoffs [31]. We found the reward as defined in Adversarial Inverse Reinforcement Learning (AIRL) [21] to be most stable when combined with the value function when applied to infinite-horizon MPPI. In the state-only setting, the policy objective becomes $r(s,s^{\prime})=\log(D(s,s^{\prime}))-\log(1-D(s,s^{\prime}))$ and the discriminator $D(s,s^{\prime})=f_{\theta}(s,s^{\prime})$. Simply put, the reward is just the logit of the discriminator $r(s,s^{\prime})=f_{\theta}(s,s^{\prime})$.
In summary, our choice of $\pi_{\text{MPPI}}$ and $r(s,s^{\prime})$ yields the MPAIL framework. As proven in Section B.2, MPAIL is indeed an AIL algorithm in the sense that it minimizes divergence from the expert policy. The procedure (Algorithm 1) itself closely resembles the original GAIL procedure. However, upon updating the value network, MPAIL does not require a policy update thereafter. We also find in practice that a temperature decay can be helpful for preventing local minima, especially in the case of online model learning. We leave a theoretical justification for this choice for future work. An overview of the training procedure can be found in Figure 1. A discussion of further meaningful implementation details, like spectral normalization, can be found in Appendix C. Though, these modifications are kept to a minimum towards our analysis of $\pi_{\text{MPPI}}$ in AIL.
@@
1:Expert state-transitions $\mathcal{D}_{E}=\{(s,s^{\prime})\}$
2:Maximum-Entropy Planner $\pi_{\text{MPPI}}$, Discriminator $D_{\theta}$, Value $V_{\phi}$
3:while not converged do
4:  Collect transitions $\left(s,s^{\prime},r_{\theta}(\cdot)\right)\in d^{\pi}$ by running $\pi_{\text{MPPI}}$ in the environment
5:  Update Discriminator parameters using $\theta$:
$\nabla_{\theta}\mathbb{E}_{s,s^{\prime}\sim d^{\pi}}[log(D_{\theta}(s,s^{\prime}))]+\nabla_{\theta}\mathbb{E}_{s,s^{\prime}\sim d^{\pi_{E}}}[log(1-D_{\theta}(s,s^{\prime}))]$ (5)
6:  Update Value parameters $\phi$ using estimated returns:
$\nabla_{\phi}\mathbb{E}_{s\sim d^{\pi}}[(G_{t}-V_{\phi}(s))^{2}]$ (6)
7:end while

### v2

# arXiv:2507.21533v2[cs.RO] 02 Mar 2026

@@ Abstract
Humans can often perform a new task after observing a few demonstrations by inferring the underlying intent. For robots, recovering the intent of the demonstrator through a learned reward function can enable more efficient, interpretable, and robust imitation through planning. A common paradigm for learning how to plan-from-demonstration involves first solving for a reward via Inverse Reinforcement Learning (IRL) and then deploying it via Model Predictive Control (MPC). In this work, we unify these two procedures by introducing planning-based Adversarial Imitation Learning, which simultaneously learns a reward and improves a planning-based agent through experience while using observation-only demonstrations. We study advantages of planning-based AIL in generalization, interpretability, robustness, and sample efficiency through experiments in simulated control tasks and real-world navigation from few or single observation-only demonstration.
@@ 1 Introduction
Figure 1: Model Predictive Adversarial Imitation Learning (MPAIL) learns costs for a planning-based, Model Predictive Control (MPC) agent from observation-only demonstration. Interactions with these costs are simultaneously used to learn a value function for experience-based reasoning beyond the horizon of the planner.
Inverse Reinforcement Learning (IRL) offers a principled approach to imitation learning by inferring the underlying intent, or reward function, that explains expert behavior. A fundamental advantage of IRL is that this reward is often readily generalizable beyond the support of the demonstration data, enabling the discovery of new policies through interaction and plans through self-prediction. Especially when demonstrations are sparse, ambiguous, or suboptimal, IRL’s interpretability is particularly well-suited for domains where understanding preferences and ensuring reliable planning are essential, such as routing on Google Maps (Barnes et al., 2023), socially aware navigation (Kretzschmar et al., 2016), and autonomous driving (Bronstein et al., 2022).
For real-time systems, learned IRL and Inverse Optimal Control (IOC) rewards are typically deployed via Model Predictive Control (MPC) (Lee et al., 2022b; Rosbach et al., 2019; Triest et al., 2023; Das et al., 2021; Lee et al., 2021; Kuderer et al., 2015; Lee et al., 2022a). Here, the offline IRL algorithm iteratively solves a Reinforcement Learning (RL) problem in an inner loop, guided by the current reward estimate. An outer loop then updates this reward to minimize the discrepancy between the agent’s and the expert’s behavior. Once training is complete, the resulting reward is integrated with MPC for real-time planning and control.
Adversarial Imitation Learning (AIL) has made significant improvements over IRL in algorithmic complexity and sample efficiency (Ho and Ermon, 2016; Baram et al., 2017). However, the reliance on an RL policy in AIL methods complicates their use in applications with safety constraints (Lee et al., 2022b; Triest et al., 2023; Das et al., 2021). Further limited by partial observability, these deployments will often prioritize planning using a model for the sake of real-time performance, trustworthiness, and interpretability (Han et al., 2024a; Katrakazas et al., 2015; Choudhury et al., 2018).
In this work, we derive planning-based AIL, yielding key benefits:
@@
    Planning-from-Observation (PfO). Towards interpretable yet scalable imitation learning, a predictive model precludes the need for expert action data and enables access to the agent’s optimization landscape. This grants crucial insight and steerability into the agent’s decision making process even as it learns from ambiguous expert data. We further show that this improves on out-of-distribution generalization, robustness, and sample efficiency when compared to policy-based AIL. We also demonstrate how policy-based AIL is fundamentally limited by the absence of reward deployment.
@@
    Unification of IRL and MPC. Otherwise considered independent training and deployment procedures, planning-based AIL allows for end-to-end interactive learning of the entire planner. Critical online settings (e.g., dynamics, preferences, control constraints) can thus be brought into training while enabling experience-based reasoning beyond the planning horizon, which we demonstrate in this work. We also find this induces a more effective adversarial dynamic than policy-based generators when learning from partial observations in the real world.
To our knowledge, this work presents the first end-to-end planning-from-observation (PfO) framework, extending PfO to continuous spaces and interactive learning. By choosing Model Predictive Path Integral control (MPPI) (Williams et al., 2017) as the embedded planner, we further gain theoretical perspective on planning-based AIL and its relationship to the seminal GAIL objective (Ho and Ermon, 2016; Torabi et al., 2019a). Thus, we name this learning algorithm: Model Predictive Adversarial Imitation Learning (MPAIL {impale}).
@@ 2 Related Work
IRL-MPC. High-dimensional continuous control applications often require an online planner for real-time control, trustworthiness, safety, or additional constraints. When using IRL to learn a reward, online deployments of these reward functions rely on an independent online MPC procedure. To enable learning local costmaps across perception and control for off-road navigation, Lee et al. (2022b) and Triest et al. (2023) similarly propose solving the forward RL problem using MPPI but deploy the reward on a different configuration more suitable for real-time planning and control. This reward deployment framework of IRL-then-MPC is currently the dominant approach for planning in high-dimensional continuous control tasks from demonstration (Lee et al., 2022b; Rosbach et al., 2019; Triest et al., 2023; Das et al., 2021; Lee et al., 2021; Kuderer et al., 2015; Lee et al., 2022a).
Model-Based IRL and Planning-Based RL. Various other works have also explored model-based AIL (Baram et al., 2016; Bronstein et al., 2022; Sun et al., 2021). However, scope is directed at training stabilization and policy optimization rather than examining planning with the learned reward. When the reward is known, as in RL, planning-based algorithms have demonstrated considerable improvement in simulation benchmarks over existing state-of-the-art RL algorithms through developments such as: online trajectory optimization, value bootstrapping, latent state planning, policy-like or learned sampling priors, and much more (Hansen et al., 2024; Bhardwaj et al., 2021; Lowrey et al., 2019; Jawale et al., 2024). This work’s implementation of MPAIL performs online trajectory optimization and value bootstrapping. This work does not implement latent state planning nor a policy-based prior to better isolate our investigations in interpretability and planning (Fu et al., 2018; Sun et al., 2021). Other existing integrations of planning-based RL with imitation learning also rely on access to expert actions (Li et al., 2025; Yin et al., 2022). No aforementioned works, save (Jawale et al., 2024), evaluate planning-based RL or AIL on a real world platform. However, we find that it is precisely real-world and out-of-distribution settings in which planning and control is most advantageous.
@@ 3.1 The POMDP Setting and the Model Predictive Agent
We adopt the Partially Observable Markov Decision Process (POMDP) to best consider highly desirable applications of IRL in which partial observability and model-based planning play crucial roles. In an unknown world state $s_{w}\in\mathcal{S}_{w}$, the agent makes an observation $o\sim p(o|s_{w})$. From a history of observations $\mathbf{o}_{0:t}$, the Agent perceives its state $s_{t}\sim p(s|\mathbf{o}_{0:t})$. Actions $a\in\mathcal{A}$ and states $s\in\mathcal{S}$ together allow the agent to self-predict forward in time using its predictive model $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. Note that these definitions crucially suggest the partial observability of $s$ due to the implicit dependency on the observation history $\mathbf{o}_{0:t}$ through the agent’s perception (e.g. mapping (Jung et al., 2024)). However, partial observations in $\mathcal{S}$ are desirably used for demonstrations to perform IRL and AIL, as full observation history would quickly become intractable (Triest et al., 2023).
The planner itself is a model predictive agent. It is capable of performing model rollouts $\tau^{(H)}_{t}=\{s_{t^{\prime}},a_{t^{\prime}}\}_{t^{\prime}=t}^{t+H}$ such that $s_{t^{\prime}+1}=f(s_{t^{\prime}},a_{t^{\prime}})$, and each rollout can thus be evaluated with a cost $C(\tau_{t})$. The agent’s objective is to create an $H$-step action sequence $\mathbf{a}_{t:t+H}$, or plan, that best minimizes the plan’s corresponding trajectory cost. The true costs or rewards under which the expert is acting is not known. Towards learning-from-observation (LfO) and lower-level policies, we consider expert data in which only states are available, as actions may be challenging or impossible to observe (Torabi et al., 2019b).
@@ 3.2 Adversarial Imitation Learning from Observation
IRL algorithms aim to learn a cost function that minimizes the cost of expert trajectories while maximizing the cost of trajectories induced by other policies (Torabi et al., 2019a; Ho and Ermon, 2016). As the problem is ill-posed and many costs can correspond to a given set of demonstrations, the principle of maximum entropy is imposed to obtain a uniquely optimal cost. It can be shown that the entropy maximizing distribution is a Boltzmann distribution (Ziebart et al., ). The state-only IRL from observation problem can be formulated by costing state-transitions $c(s,s^{\prime})$ rather than state-actions $c(s,a)$ as in (Torabi et al., 2019a):
$\text{IRLfO}_{\psi}(\pi_{E})=\operatorname*{arg\,max}_{c\in\mathbb{R}^{\mathcal{S}\times\mathcal{S}}}-\psi(c)+\left(\min_{\pi\in\Pi}-\lambda\mathbb{H}(\pi)+\mathbb{E}_{\pi}[c(s,s^{\prime})]\right)-\mathbb{E}_{\pi_{E}}[c(s,s^{\prime})],$ (1)
@@
As shown in (Ho and Ermon, 2016; Torabi et al., 2019a), this objective can be shown to be dual to the Adversarial Imitation Learning (AIL) objective under a specific choice of cost regularizer $\psi$,
@@
where $D(\cdot)$ is the discriminator function. The exact form of $D$ has consequences on the policy objective and differs by AIL algorithm. Now equipped with the optimization objective, we continue in our derivation of planning-based AIL by choosing the form of the policy class and reward function.
@@ 3.3 Choosing a Policy and Reward
To reiterate, we set our sights on the AIL objective (Equation 2) which aims to simultaneously learn reward and policy from demonstration. However, the formulation remains intimately connected with policy optimization through the assumption of an RL procedure. Towards planning-based optimization, we proceed by modifying the RL formulation in Equation 1 as described in (Ho and Ermon, 2016; Fu et al., 2018). Similar to (Bhardwaj et al., ), we replace the entropy loss $\lambda\mathbb{H}(\pi)$ with a Kullback-Leibeler (KL) divergence constraint on the previous policy $\overline{\pi}$:
@@
We then observe that a choice of planner satisfies the RL objective as in Equation 3. By choosing Model Predictive Path Integral (MPPI) as the planner, as proven in Section B.1, we solve an equivalent problem provided the MDP is uniformly ergodic. Namely, MPPI solves for a KL-constrained cost-minimizer over trajectories (Bhardwaj et al., ):
@@
1:Expert state-transitions $\mathcal{D}_{E}=\{(s,s^{\prime})\}$ Maximum-Entropy Planner $\pi_{\text{MPPI}}$, Discriminator $D_{\theta}$, Value $V_{\phi}$
2:while not converged do
3:  Collect transitions $\left(s,s^{\prime},r_{\theta}(\cdot)\right)\in d^{\pi}$ by running $\pi_{\text{MPPI}}$ (Alg. 3) in the environment
4:  Update Discriminator parameters $\theta$ with Binary Cross Entropy (BCE) loss:
$\nabla_{\theta}\mathbb{E}_{s,s^{\prime}\sim d^{\pi}}[\log(D_{\theta}(s,s^{\prime}))]+\nabla_{\theta}\mathbb{E}_{s,s^{\prime}\sim d^{\pi_{E}}}[\log(1-D_{\theta}(s,s^{\prime}))]$ (5)
5:  Update Value parameters $\phi$ using estimated returns:
$\nabla_{\phi}\mathbb{E}_{s\sim d^{\pi}}[\left(G_{t}-V_{\phi}(s)\right)^{2}]$ (6)
6:end while
Model rollouts must practically be limited to some timestep length $H$, resulting in myopic plans and limiting applications to short-horizon tasks (Bhardwaj et al., 2021). To resolve this, infinite-horizon MPPI bootstraps the final states in the rollout using a terminal cost-to-go, or value, function (Bhardwaj et al., 2021; Hatch and Boots, 2021; Lowrey et al., 2019) $V_{\phi}:\mathcal{S}\rightarrow\mathbb{R}$ that estimates the expected return $G_{t}$ of a state $s_{t}$ as $G_{t}=\mathbb{E}_{\pi}[R_{t+1}+\gamma R_{t+2}+...|S_{t}=s_{t}]$, where $R_{t+1}=R(s_{t},s_{t+1})$. The result of MPPI’s approximately global policy optimization at each timestep is what is referred to as the MPPI policy, $\pi_{\text{MPPI}}$. Section B.1 proves how this formulation can be equivalent to the entropy-regularized RL objective, while also in the observation-only setting. Pseudocode for the MPPI procedure can be found in Algorithm 2 as well as for its adaptation as an RL policy in Algorithm 3. Figure 2 illustrates the policy.
Figure 2: Illustration of $\pi_{\text{MPPI}}$ in MPAIL. (1) A set of action sequences (plans) are sampled and rolled out. (2) Plans are costed according to the discriminator, shifting the distribution towards the expert. Temperature $\lambda$ optionally decays over episodes, narrowing the distribution. (3) The policy $\pi_{\text{MPPI}}$ is the result of a Gaussian fit to the optimized plans and their respective first actions.
Our chosen AIL agent, infinite-horizon MPPI, can be viewed as optimizing for a new policy at every state. As in the online learning perspective (Wagener et al., 2019), these MPPI optimizations can occur online rather than offline as part of, for instance, a slower actor-critic update. By deconstructing the agent this way, we require not the policy to generalize but the reward.
Provided the agent, we now proceed with selecting its objective. Recent work has shown many potential choices of valid policy objectives, each with various empirical tradeoffs (Orsini et al., 2021). We found the reward as defined in Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2018) to be most stable when combined with the value function when applied to infinite-horizon MPPI. In the state-only setting, the policy objective becomes $r(s,s^{\prime})=\log(D(s,s^{\prime}))-\log(1-D(s,s^{\prime}))$ and the discriminator $D(s,s^{\prime})=\sigma\circ d_{\theta}(s,s^{\prime})$. Simply put, the reward is the logit of the discriminator $r(s,s^{\prime})=d_{\theta}(s,s^{\prime})$.
In short, our choice of $\pi_{\text{MPPI}}$ and $r(s,s^{\prime})$ yields MPAIL. As proven in Section B.2, MPAIL is indeed an AIL algorithm in the sense that it minimizes divergence from the expert policy. The procedure (Algorithm 1) itself closely resembles the original GAIL procedure. However, upon updating the value network, MPAIL does not require a policy update thereafter, since “policies” are in theory solved online. In practice a temperature decay can be helpful for preventing early collapse, especially in the case of online model learning. We leave a theoretical justification for this choice for future work. An overview of the training procedure can be found in Figure 1. A discussion of further meaningful implementation details, like spectral normalization, can be found in Appendix C. Though, these modifications are kept to a minimum towards our analysis of $\pi_{\text{MPPI}}$ in AIL.